### v2

# arXiv:2602.23010v2[cs.GR] 01 Mar 2026

Abstract
We present Helmlab, a 72-parameter analytical color space for UI design systems. The forward transform maps CIE XYZ to a perceptually-organized Lab representation through learned matrices, per-channel power compression, Fourier hue correction, and embedded Helmholtz–Kohlrausch lightness adjustment. A post-pipeline neutral correction guarantees that achromatic colors map to a=b=0a\!=\!b\!=\!0 (chroma $10^{-6}$), and a rigid rotation of the chromatic plane improves hue-angle alignment without affecting the distance metric, which is invariant under isometries. On the combvd dataset (3,813 color pairs), Helmlab achieves a stress of 23.30, a 20.2% reduction from CIEDE2000 (29.18). A blue-band refit with sub-dataset penalties reduces gradient non-uniformity in the blue–cyan region by $8.9\times$ at a cost of only +0.08+0.08 stress. Cross-validation on He et al. 2022 and MacAdam 1974 shows competitive cross-dataset performance. The transform is invertible with round-trip errors below $10^{-14}$. Gamut mapping, design-token export, and dark/light mode adaptation utilities are included for use in web and mobile design systems.
Keywords: color space, color difference, perceptual uniformity, STRESS, Helmholtz–Kohlrausch, UI design systems
1 Introduction
UI design systems need accurate perceptual distance prediction for contrast checking, an achromatic axis where grays map to zero chroma for gradient interpolation, and reasonable hue-angle alignment for programmatic color manipulation. No existing color space provides all three.
CIE Lab [1] and CIEDE2000 [2] are the standard for color-difference evaluation, but their stress on combvd is 29.18. Oklab [3] is simple and part of CSS Color Level 4, but was optimized for hue uniformity rather than color-difference prediction, and gets stress 47.5 (Euclidean) on combvd. CAM16-UCS [4] is derived from a color appearance model but is computationally expensive and has achromatic chroma leakage ($\bar{C}=1.42$). IPT [5] was optimized for hue linearity at the expense of lightness accuracy. $J_{z}a_{z}b_{z}$ [6] targets HDR via the PQ curve but is complex and not CSS-native. None of these spaces jointly optimize the color-space transform and distance metric end-to-end against psychophysical data.
In this paper, we treat the entire XYZ-to-distance pipeline as a single system with 72 learnable parameters and optimize all of them jointly against the largest available color-difference dataset. The contributions are:
  1. 1.
    A 72-parameter analytical color space with stress 23.30 on combvd, 20.2% lower than CIEDE2000.
  2. 2.
    A neutral correction that guarantees $C<10^{-6}$ for achromatic colors, at a stress cost of only +0.04+0.04.
  3. 3.
    A rigid rotation in the $ab$ plane that reduces hue error (RMS 18.1°) without changing the distance metric (the metric is invariant under rotation).
  4. 4.
    Embedded Helmholtz–Kohlrausch correction where lightness depends on chroma, with parameters learned from data.
  5. 5.
    A blue-band refit that reduces gradient non-uniformity in the blue–cyan region by $8.9\times$ while improving stress on 5 of 6 combvd sub-datasets.
  6. 6.
    Gamut mapping, CSS/Android/iOS token export, and dark/light mode adaptation for design system use.
2 Design Goals and Metrics
Helmlab is optimized for UI design systems operating in sRGB and Display P3 gamuts under typical viewing conditions (D65 illuminant, average surround). We define five quantitative targets:
Color-difference accuracy (stress).
The STRESS metric [7] measures agreement between predicted distances $\Delta E_{i}$ and visual differences $DV_{i}$:
$\mathrm{STRESS}=100\sqrt{\frac{\sum_{i}(DV_{i}-F\cdot\Delta E_{i})^{2}}{\sum_{i}DV_{i}^{2}}}$ (1)
where $F=\sum DV_{i}\Delta E_{i}/\sum\Delta E_{i}^{2}$. Lower is better. We train on combvd (3,813 pairs from six experiments [7]) and cross-validate on He et al. 2022 [8] (82 pairs) and MacAdam 1974 [9] (128 pairs).
Achromatic integrity.
Neutral colors (D65-adapted grays, i.e.,  stimuli on the achromatic axis proportional to the D65 white point) must map to $a=b=0$. We measure maximum chroma $C_{\max}$ over 256 gray levels spanning $Y\in[0.001,2.0]$.
Hue-angle alignment.
[…]
Round-trip accuracy.
For any sRGB color, $\|T^{-1}(T(\mathbf{x}))-\mathbf{x}\|_{\infty}$ should be below $10^{-8}$.
Munsell uniformity.
Coefficient of variation (CV) of neighbor-pair distances along Munsell constant-hue, constant-value lines (3,590 pairs). Lower CV indicates more uniform spacing.
3 The Helmlab Color Space
3.1 Forward Transform
The forward transform maps CIE XYZ (D65) to Lab through eleven stages (Figure˜1). All parameters are jointly optimized (Section˜5).
Figure 1: The Helmlab forward transform pipeline. Blue: linear operations; yellow: nonlinear corrections; green: structural guarantees (NC, rotation). All 72 parameters are jointly optimized.
Stage 1: Cone response (9 params).
A learned $3\!\times\!3$ matrix $\mathbf{M}_{1}$ maps XYZ to cone-response-like signals:
$\mathbf{c}=\mathbf{M}_{1}\begin{bmatrix}X\\ Y\\ Z\end{bmatrix}$ (2)
Unlike fixed matrices (Hunt–Pointer–Estévez, Bradford), $\mathbf{M}_{1}$ is freely optimized.
Stage 2: Power compression (3 params).
Each channel undergoes signed power compression:
$c_{i}^{\prime}=\mathrm{sign}(c_{i})\,|c_{i}|^{\gamma_{i}},\quad i\in\{0,1,2\}$ (3)
The exponents $\gamma_{i}$ generalize the cube-root of CIE Lab ($\gamma=1/3$). The optimized values are close to 0.4, which is between the cube-root (0.33) and square-root (0.5) compressions used in other spaces.
Stage 3: Lab projection (9 params).
A second matrix $\mathbf{M}_{2}$ projects compressed signals to opponent channels:
$\begin{bmatrix}L_{\mathrm{raw}}\\ a_{\mathrm{raw}}\\ b_{\mathrm{raw}}\end{bmatrix}=\mathbf{M}_{2}\,\mathbf{c}^{\prime}$ (4)
Stage 4: Hue correction (8 params).
A 4-harmonic Fourier rotation corrects systematic hue distortions:
$\delta(h)=\sum_{k=1}^{4}\!\bigl[\alpha_{k}\cos(kh)+\beta_{k}\sin(kh)\bigr]$ (5)
where $h=\mathrm{atan2}(b,a)$. The chromatic vector $(a,b)$ is rotated by $\delta(h)$, preserving chroma while redistributing hue angles.
Stage 5: Helmholtz–Kohlrausch embedding (6 params).
Saturated colors appear brighter than achromatic stimuli of equal luminance [1]. We embed this effect in the lightness channel directly:
$L\mathrel{+}=w_{\mathrm{HK}}\cdot C^{p_{\mathrm{HK}}}\cdot\bigl[1+f_{\mathrm{HK}}(h)\bigr]$ (6)
where $C=\sqrt{a^{2}+b^{2}}$ and $f_{\mathrm{HK}}(h)$ is a 2-harmonic hue modulation. The optimized weight is $w_{\mathrm{HK}}=0.389$ and the power is $p_{\mathrm{HK}}=0.849$, making lightness explicitly dependent on chroma. Most existing spaces treat lightness and chroma as independent.
Stage 6: Lightness refinement (8 params).
A cubic polynomial (3 params) with hue-dependent additive term (2 params), followed by dark-region exponential compression (3 params):
$L_{1}$ $=p_{1}L^{3}+p_{2}L^{2}+p_{3}L+t\cdot f_{\mathrm{Lh}}(h)$ (7)
$L_{2}$ $=L_{1}\cdot\exp\!\bigl[\lambda_{d}(h)\cdot L_{1}(1-L_{1})^{2}\bigr]$ (8)
where t=L(1L)t=L(1-L) and $f_{\mathrm{Lh}}(h)$ is a 1-harmonic Fourier correction targeting the hue–lightness interaction sign flip (dark cyans vs mid cyans). The dark compression coefficient $\lambda_{d}$ includes hue modulation to handle the known difference in dark-blue vs dark-yellow perception.
Stages 7–8: Chroma processing (18 params).
Chroma processing proceeds in four interleaved sub-stages: (a) hue-dependent scaling (4 harmonics, 8 params):
$(a,b)\mathrel{*}=\exp\!\Bigl[\sum_{k=1}^{4}\!\bigl[\alpha^{\prime}_{k}\cos(kh)+\beta^{\prime}_{k}\sin(kh)\bigr]\Bigr]$ (9)
[…]
Stage 9: Hue-dependent lightness (4 params).
A final $L\mathrel{*}=\exp[g(h)]$ captures residual hue–lightness interactions (e.g.,  blue appears darker, yellow lighter at equal luminance).
Stage 10: Neutral correction.
With independent $\gamma_{i}$ per channel, the D65 achromatic axis (stimuli proportional to $X_{D65},Y_{D65},Z_{D65}]$) generally does not map to a=b=0a\!=\!b\!=\!0 after the enrichment stages. We correct this by running 256 gray levels ($Y\in[0.001,2.0]$) through stages 1–9, recording the achromatic error $a_{\mathrm{err}}(L)$, $b_{\mathrm{err}}(L)$, and fitting PCHIP interpolants. At runtime:
$a\leftarrow a-a_{\mathrm{err}}(L),\quad b\leftarrow b-b_{\mathrm{err}}(L)$ (10)
Because the correction is applied after all nonlinear stages, it accounts for every distortion in the pipeline. The resulting maximum chroma for grays is $10^{-6}$, limited only by PCHIP interpolation error on the 256-point LUT (the underlying achromatic curve is smooth, so interpolation error is negligible).
Stage 11: Rigid rotation ($\varphi$).
A uniform rotation of the $(a,b)$ plane by angle $\varphi$:
$\begin{bmatrix}a^{\prime}\\ b^{\prime}\end{bmatrix}=\begin{bmatrix}\cos\varphi&-\sin\varphi\\ \sin\varphi&\cos\varphi\end{bmatrix}\begin{bmatrix}a\\ b\end{bmatrix}$ (11)
This is an isometry: it preserves $\Delta a^{2}+\Delta b^{2}$ for any pair, and therefore preserves the distance metric exactly (Section˜3.2). The rotation angle $\varphi=-28.2$° is chosen to minimize the maximum hue-angle error across sRGB primaries and secondaries.
3.2 Rotation Invariance
The distance metric (Section˜4) depends on $\Delta L$, $\Delta a^{2}+\Delta b^{2}$, $\bar{L}$, and $\bar{C}=\sqrt{\bar{a}^{2}+\bar{b}^{2}}$. A rigid rotation of the $(a,b)$ plane preserves each of these quantities: $\Delta a^{\prime 2}+\Delta b^{\prime 2}=\Delta a^{2}+\Delta b^{2}$ and $\bar{a}^{\prime 2}+\bar{b}^{\prime 2}=\bar{a}^{2}+\bar{b}^{2}$. Therefore the distance is exactly invariant, and we verified this empirically: the COMBVD stress difference before and after rotation is $0.0000000000$.
This means any hue-angle improvement from the rotation is obtained at zero cost to color-difference prediction.
3.3 Surround Parameter
Helmlab includes provisions for a surround luminance parameter $S\in[0,1]$ (0=dark, 0.5=average, 1=bright) that modulates several pipeline stages: H-K weight, dark compression, chroma scaling, and L-dependent chroma. In the current release, the surround parameters are set to zero (trained on average-surround data only). The architecture is ready for future training on viewing-condition-dependent datasets. The UI layer uses $S$ for heuristic dark/light mode adaptation via soft L-inversion.
3.4 Inverse Transform
Each stage has an explicit inverse. Matrix stages invert via $\mathbf{M}_{1}^{-1}$, $\mathbf{M}_{2}^{-1}$; power compression via $c_{i}=\mathrm{sign}(c_{i}^{\prime})|c_{i}^{\prime}|^{1/\gamma_{i}}$ (closed form). Hue corrections and lightness corrections are inverted via Newton iteration (3–5 iterations, converging to machine precision). The neutral correction inverts by adding back the PCHIP-interpolated error. Round-trip accuracy across the full sRGB gamut is $10^{-14}$.
4 Distance Metric
Given two colors with coordinates $(L_{1},a_{1},b_{1})$ and $(L_{2},a_{2},b_{2})$, we compute:
$S_{L}$ $=1+s_{L}(\bar{L}-0.5)^{2}$ (12)
$S_{C}$ $=1+s_{C}\bar{C}$ (13)
where $\bar{L}$, $\bar{C}$ are pair averages. The raw distance is:
$d=\left[\left(\frac{\Delta L}{S_{L}}\right)^{\!2}+w_{C}\frac{\Delta a^{2}+\Delta b^{2}}{S_{C}^{2}}\right]^{p/2}$ (14)
followed by monotonic compression and post-power:
$\Delta E=\left[\frac{d}{1+c\cdot d}\right]^{q}$ (15)
The pair-dependent weights $S_{L}$ and $S_{C}$ are inspired by CIEDE2000 but learned from data (optimized values: $s_{L}=0.001$, $s_{C}=0.022$). The Minkowski exponent $p=0.804$ and chroma weight $w_{C}=1.046$ slightly favor chroma differences over lightness. The compression $c=1.590$ and post-power $q=1.1$ reduce systematic bias for large color differences.
The distance metric contributes 7 parameters. Total system: 9+3+9+8+6+8+18+4 = 65 space + 7 metric = 72 parameters.
5 Optimization
All 72 parameters are jointly optimized by minimizing:
$\mathcal{L}=\mathrm{STRESS}_{\textsc{combvd}}+0.05\cdot\mathrm{STRESS}_{\mathrm{He}}+\mathcal{L}_{\mathrm{blue}}$ (16)
using L-BFGS-B [10] with box constraints. We perform 10 random restarts (each 5,000 iterations) and keep the best. The He et al. term ($\lambda=0.05$) acts as a mild regularizer against overfitting to combvd.
The blue-band penalty is:
$\mathcal{L}_{\mathrm{blue}}=\lambda_{b}(\mathrm{STRESS}_{\mathrm{RIT}}+\mathrm{STRESS}_{\mathrm{LEEDS}}+0.5\,\mathrm{STRESS}_{\mathrm{FullBlue}})$ (17)
where $\lambda_{b}=0.1$ and the subscripts denote combvd sub-datasets with high blue-band content. Additional constraints enforce round-trip accuracy ($\lambda_{\mathrm{RT}}=100$), achromatic integrity ($\lambda_{\mathrm{ach}}=500$), and Munsell uniformity ($0.1\cdot\max(\mathrm{CV}-20,0)$). The blue-band penalty was introduced in v2 of this paper to address gradient non-uniformity in the blue–cyan region (Section˜6.6).
Training data.
combvd comprises 3,813 color pairs from six psychophysical experiments: BFD-P(D65) (2,028), BFD-P(M) (548), BFD-P(C) (200), Witt (418), RIT-DuPont (312), Leeds (307). All pairs have visual-difference values on a common scale. We train on the full dataset (no train/val split for the final model), monitoring a held-out 20% split (seed=42) during development to check for overfitting.
Cross-validation data.
He et al. 2022 (82 pairs, 10° observer) enters the loss as a regularizer ($\lambda\!=\!0.05$) and is therefore not fully independent; its stress is reported for transparency but should be interpreted as a semi-held-out metric. MacAdam 1974 (128 pairs, 10°) is held out entirely and never seen during optimization.
Convergence.
The optimizer typically converges within 3,000–5,000 function evaluations per restart. Five-fold cross-validation (with re-optimization per fold) yields a mean validation gap of +0.98+0.98 stress points (std. $0.78$), indicating mild but acceptable overfitting for a 72-parameter model.
6 Evaluation
6.1 Color-Difference Prediction
Figure˜2 and Table˜1 show stress on combvd for Helmlab and nine baseline methods. Helmlab achieves 23.30, compared to 29.18 for CIEDE2000 (-20.2%). A 10,000-iteration paired bootstrap gives Helmlab 95% CI $[22.50,23.93]$ and CIEDE2000 95% CI $[27.64,30.84]$. The intervals do not overlap ($p<10^{-4}$).
Figure 2: stress on COMBVD (3,813 pairs). Lower is better.
Table 1: stress on combvd (3,813 pairs). Lower is better.
Method stress vs CIEDE2000
Helmlab 23.30 -20.2%
CIEDE2000 29.18
$\Delta E^{*}_{ab}$ (CIE76) 42.80 +46.7%
CIE94 33.59 +15.1%
$\Delta E$ CMC 34.04 +16.6%
CAM16-UCS (Euclid.) 33.90 +16.2%
IPT (Euclid.) 41.21 +41.3%
$J_{z}a_{z}b_{z}$ (Euclid.) 41.90 +43.6%
Oklab (Euclid.) 47.46 +62.7%
sRGB (Euclid.) 67.82 +132.5%
Figure 3: Predicted vs observed color differences on COMBVD.
The baselines in Table˜1 use Euclidean distance in each space’s native Cartesian coordinates, without parametric weighting or compression. Oklab’s high stress (47.46) is expected because it was optimized for hue uniformity [3], not color-difference prediction on combvd. The Helmlab distance metric includes pair-dependent weighting ($S_{L}$, $S_{C}$), a Minkowski exponent, and monotonic compression, so the comparison is not on equal terms. With Euclidean distance only, Helmlab gets stress 27.6, which already improves on CIEDE2000 (29.18) by 5.4%. The full distance metric then reduces stress by a further 4.3 points. Relative to CIEDE2000, the space transform and the distance metric contribute roughly equally to the total improvement.
6.2 Cross-Validation
On He et al. 2022 (regularizer, 82 pairs), Helmlab achieves stress 30.3 vs CIEDE2000’s 32.6. On MacAdam 1974 (held out entirely, 128 pairs), Helmlab achieves 20.3 vs CAM16-UCS’s 18.7 (best). Figure˜4 visualizes these results. Helmlab outperforms CIEDE2000 on He et al. but trails CAM16-UCS on MacAdam. This is consistent with what one would expect from a model trained primarily on combvd: it generalizes reasonably but does not dominate on all evaluation sets.
Figure 4: Cross-dataset validation on held-out data.
6.3 Generation Properties: The Measurement–Generation Tradeoff
A color space optimized for distance prediction does not necessarily produce usable coordinates for color generation. During development, the measurement-optimal configuration mapped neutral grays to chroma $\bar{C}\approx 0.34$. The optimizer had deformed the achromatic axis to reduce stress, which made gray gradients unusable because they showed visible color artifacts. Adding a penalty to constrain grays to $C=0$ did not help: the optimizer either stayed in the measurement-optimal basin or escaped to a different basin with much worse stress (26–31). There was no smooth transition between the two.
The neutral correction (Section˜3, Stage 10) addresses this problem. It subtracts the achromatic error after all enrichment stages, so the optimizer can find its measurement-optimal parameters without constraint. The achromatic axis is then corrected separately. The stress increase from the neutral correction alone is +0.04+0.04 (from a measurement-only baseline of 23.18 to 23.22); the subsequent blue-band refit (Section˜6.6) adds another +0.08+0.08, yielding the final 23.30. Achromatic chroma drops from 0.34 to $10^{-6}$.
Figure˜5 shows neutral ramp uniformity and achromatic chroma leakage. Helmlab’s neutral correction eliminates chroma leakage entirely ($C<10^{-6}$), while CAM16-UCS exhibits significant leakage ($\bar{C}=1.42$). Table˜2 summarizes all generation properties.
Figure 5: Left: neutral ramp step uniformity. Right: achromatic chroma leakage. Helmlab’s NC guarantees $C<10^{-6}$ for all grays.
Table 2: Generation quality metrics.
Property Value
Achromatic max $C$ $10^{-6}$
Hue RMS (6 primaries) 18.1°
Hue max error 23.9°
Round-trip $\|\cdot\|_{\infty}$ $10^{-14}$
Munsell CV 20.0%
Jacobian det (min) 0.10
Jacobian cond (median) 7.7
5-fold CV gap (mean) +0.98
He stress 30.33
MacAdam stress 20.33
The achromatic guarantee means gray gradients and lightness ramps interpolate without color artifacts, which is necessary for design-token generation and palette export. The hue RMS of 18.1° means sRGB primaries land within about 24° of their expected positions. This is acceptable for palette generation but not perfect. The rigid rotation (Section˜3, Stage 11) provides this hue improvement without affecting the distance metric because the metric is invariant under rotation. The Jacobian was evaluated on a $64\!\times\!64\!\times\!64$ sRGB grid via finite differences; the minimum determinant (0.10) is positive throughout, so the map is locally invertible and orientation-preserving across the entire sRGB gamut.
6.4 Ablation Study
Table˜3 isolates the contribution of key components using frozen ablation: each row disables one component while keeping all other parameters at their optimized values. This measures the marginal contribution of each component given the current parameter setting, not its isolated effect.
Table 3: Ablation: effect of removing components (frozen, not re-optimized).
Configuration stress $\Delta$
Full Helmlab 23.30
Euclidean distance only 27.60 +4.3
No H-K embedding ($w_{\mathrm{HK}}=0$) 27.30 +4.0
No hue correction 38.81 +15.5
No dark L compression 23.49 +0.2
Rotation (-28.2°) 23.30 0.00
No rotation 23.30 0.00
With Euclidean distance, the space alone gets stress 27.6, which already improves on CIEDE2000 (29.18) by 5.4%. The full metric brings it down to 23.30, recovering a further 4.3 points. Compared to other Euclidean baselines (CAM16-UCS = 33.9), the space alone gives a substantial improvement. Among individual stages, hue correction is the most critical ($\Delta=+15.5$), followed by the H-K embedding ($\Delta=+4.0$). These large frozen-ablation deltas reflect the co-optimization of all parameters: disabling one stage disrupts the balance of the entire pipeline. The rotation row confirms that the metric is exactly invariant under rotation.
6.5 Sub-Dataset Performance
Table˜4 breaks down stress by combvd sub-dataset. Helmlab outperforms CIEDE2000 on 3 of 6 sub-datasets (BFD-P(D65), BFD-P(M), Witt). On the blue-heavy sub-datasets (LEEDS, RIT-DuPont), the blue-band refit reduces the gap substantially compared to an earlier version without the penalty: RIT-DuPont improves from 23.49 to 20.41, nearly matching CIEDE2000’s 19.51. BFD-P(C) is the only sub-dataset measured under illuminant C (rather than D65) and with a 2° observer, which may explain the remaining deficit.
Table 4: stress on individual combvd sub-datasets.
Sub-dataset NN Helmlab CIEDE2000
BFD-P(D65) 2,028 23.11 24.14
BFD-P(M) 548 22.22 35.09
BFD-P(C) 200 30.15 30.05
Witt 418 28.88 30.37
Leeds 307 20.92 19.46
RIT-DuPont 312 20.41 19.51
6.6 Blue-Band Analysis
During development, we observed that an initial version of Helmlab (without the blue-band penalty) exhibited severe gradient non-uniformity in the blue–cyan region. A red-to-blue gradient interpolated in Helmlab Lab coordinates showed a $51.4\times$ max/min step-size ratio when measured with CIEDE2000, compared to $1.3\times$ for the same gradient in CIE Lab. Frozen ablation traced the problem to the 4-harmonic hue correction, whose Fourier coefficients overshoot in the blue–magenta band ($h\approx 240$340340°). Disabling hue correction reduced the gradient ratio from $51.4\times$ to $4.9\times$ but increased combvd stress by +15.5+15.5 points, making this an unacceptable trade-off.
The solution was to add explicit sub-dataset penalties to the loss function (Equation˜17) and re-optimize all 72 parameters. With $\lambda_{b}=0.1$, the red-to-blue gradient ratio drops from $51.4\times$ to $5.8\times$ ($8.9\times$ improvement), while combvd stress increases by only +0.08+0.08 (from 23.22 to 23.30). stress improves on 5 of 6 sub-datasets (Table˜4), with the largest gain on RIT-DuPont (3.08-3.08 points). The only regression is BFD-P(D65) (+0.49+0.49), the largest sub-dataset, which accounts for the slight overall stress increase.
7 Practical Adoption
Figure˜6 shows the sRGB gamut boundary in Helmlab coordinates at three lightness levels. The boundaries are smooth and well-behaved. Helmlab includes a Python utility layer for design system integration:
Figure 6: sRGB gamut in Helmlab aabb plane at three L levels. Points colored by their sRGB values.
Gamut mapping.
Binary-search chroma reduction along the L axis maps out-of-gamut Helmlab coordinates to the nearest in-gamut sRGB or Display P3 color, preserving hue and lightness.
Contrast utilities.
ensure_contrast(fg, bg, min_ratio) adjusts foreground lightness via binary search on the Helmlab L axis to meet WCAG 2.x contrast requirements (4.5:1 for AA, 7:1 for AAA). Hue and chroma are preserved.
Palette generation.
Lightness ramps and hue rings at fixed chroma produce perceptually-spaced palettes. Semantic scales (Tailwind-style 50–950) are generated by interpolating L between 0.97 and 0.10 with chroma/hue preserved.
Dark/light mode.
When surround parameters are trained, the surround $S$ directly modulates the transform. Currently, a fallback soft L-inversion maps colors between light ($S=0.7$) and dark ($S=0.2$) modes with contrast enforcement.
Token export.
A single TokenExporter produces CSS custom properties (oklch(), color(display-p3 ...)), Android XML (using HCT tone/chroma/hue semantics for Material Design compatibility), iOS Swift (Display P3), Tailwind config, and raw JSON.
8 Limitations and Future Work
There are several limitations worth noting:
  1. 1.
    Training data. combvd predominantly uses the CIE 1964 10° standard observer ($\sim$95% of pairs). Only BFD-P(C) (200 pairs, 5%) uses the 2° observer. Gao et al. 2023 [11] showed that the difference between 2° and 10° observers is not statistically significant for stress evaluation, and MacAdam 1974 (measured with the 2° observer) cross-validation confirms reasonable generalization.
  2. 2.
    Model complexity. 72 parameters is considerably more than Oklab ($\sim$15 designed parameters). Each stage has a perceptual motivation, but the overall model is harder to interpret than simpler spaces.
  3. 3.
    Hue alignment. RMS hue error of 18.1° is a compromise. Some hue regions (cyan) have errors up to 23.9°. Further harmonic terms could improve this at the risk of overfitting.
  4. 4.
    Surround dependency. The surround parameter $S$ is architecturally present but not yet trained on viewing-condition data. Current dark/light adaptation uses heuristics.
  5. 5.
    Sub-dataset variation. Helmlab underperforms CIEDE2000 on 3 of 6 combvd sub-datasets (Table˜4). The blue-band refit improves 5 of 6 sub-datasets compared to the previous version; only BFD-P(C), measured under illuminant C with a 2° observer, remains notably worse (30.15 vs 30.05).
  6. 6.
    Not general-purpose. Helmlab is optimized for UI workflows (sRGB/P3 gamut, screen viewing). It is not designed for print, photography, or spectral applications.
  7. 7.
    Blue-band trade-off. The blue-band refit (Section˜6.6) reduces gradient non-uniformity by $8.9\times$ at a cost of +0.08+0.08 stress on combvd. The Munsell CV increases from 17.6% to 20.0%, hue RMS from 16.1° to 18.1°, and the He et al. cross-validation stress increases from 29.0 to 30.3.
  8. 8.
    Statistical reporting. Bootstrap confidence intervals confirm the aggregate improvement over CIEDE2000 ($p<10^{-4}$), but per-sub-dataset significance and parameter uncertainty across restarts remain unquantified.
The open-source release includes a second, generation-optimized space (GenSpace) with shared $\gamma\!=\!\tfrac{1}{3}$ and no enrichment stages, which trades distance accuracy for faster, invertible gradient and palette generation. For future work, we plan to train the surround parameter on paired light/dark viewing data and to expand the training set with targeted human feedback using the bidirectional optimization framework that is already in the codebase.
9 Conclusion
We have presented Helmlab, a 72-parameter analytical color space that achieves stress 23.30 on combvd, which is 20.2% lower than CIEDE2000. The achromatic axis is guaranteed to have $C<10^{-6}$ via the neutral correction, and a rigid rotation brings hue error down to 18.1° RMS without affecting the distance metric. A blue-band refit with sub-dataset penalties reduces gradient non-uniformity in the blue–cyan region by $8.9\times$ at a cost of only +0.08+0.08 stress points, while improving stress on 5 of 6 sub-datasets. The neutral correction is the key element that makes the space usable for both measurement and generation: it lets the optimizer find its best parameters freely and then fixes the achromatic axis. Code, trained parameters, and utilities are available at https://github.com/Grkmyldz148/helmlab. An interactive demo is at https://grkmyldz148.github.io/helmlab/demo.html and documentation at https://grkmyldz148.github.io/helmlab/.
Acknowledgments
Thanks to the 71 anonymous observers who participated in the bidirectional evaluation study during development.
Appendix A Parameter Table
Table˜5 lists all 72 optimized parameters grouped by pipeline stage.
Table 5: Optimized Helmlab parameters (v20b, $\varphi=-28.2$°).
Stage Parameter Value Description
$\mathbf{M}_{1}$ (9) Row 0 [0.734, 0.240, -0.158] XYZ $\to$ cone-like
Row 1 [-0.330, 1.235, -0.000]
Row 2 [0.086, 0.341, 0.832]
$\gamma$ (3) $\gamma_{0},\gamma_{1},\gamma_{2}$ 0.389, 0.416, 0.424 Power compression
$\mathbf{M}_{2}$ (9) Row 0 [-0.420, 0.485, 0.642] Compressed $\to$ Lab
Row 1 [1.933, -2.718, 0.764]
Row 2 [0.006, 1.628, -1.261]
Hue corr. (8) $\alpha_{1..4},\beta_{1..4}$ see JSON 4-harmonic Fourier
H-K (6) $w,p,m,s_{1},c_{2},s_{2}$ 0.389, 0.849, -0.494, … Chroma $\to$ L
Light. (8) $p_{1},p_{2},p_{3},Lh_{c1},Lh_{s1},\lambda_{d},h_{c},h_{s}$ 0.222, -0.727, 0.695, … Cubic + Lh + dark
Chroma (18) 4-harm + power + $lc_{1},lc_{2}$ + HLC see JSON Hue/power/L-dep
Hue-L (4) $g_{c1..2},g_{s1..2}$ see JSON Hue $\to$ L
Distance (7) $s_{L}$ $1.01\times 10^{-3}$ Pair-dep L weight
$s_{C}$ 0.022 Pair-dep C weight
pp 0.804 Minkowski exponent
$w_{C}$ 1.046 Chroma weight
$c$ 1.590 Compression
$q$ 1.100 Post-power
$\alpha$ 0.000 Linear term
Full parameter values are available in the accompanying metric_params.json file.